Randomly polarized light of intensity i0 incident on two sheet polarizers . The transmission axis of the first polarizer is vertical that of The second makes a 30 angle with the vertical The intensity and polarization State of the light after passing through the two is....


1
Expert's answer
2021-05-27T08:31:11-0400

"I_1=I_0\sin^2 30°=\frac{I_0}{4},"

"I_2=I_1\sin^2(90°-30°)=\frac{3}{16}I_0."